A box has 2 red chips, 2blue chips, and 1 green chip. Two chips are randomly selected together. What is the probability of selecting 2 blue chips? List the elements
Answer by ewatrrr(24785)


Hi,
5 chips total originally, of which 2 are red, 2 are blue and 1 is green
Two chips are randomly selected 'together'
P (selecting 2 blue chips) = 2/5 * 1/4 = 2/20 = 1/10
10% chance of selecting two blue chips.